# Gaussian Splatting with Localized Points Management

### v2

@@ Abstract
Point management is a critical component in optimizing 3D Gaussian Splatting (3DGS) models, as the point initiation (e.g., via structure from motion) is distributionally inappropriate. Typically, the Adaptive Density Control (ADC) algorithm is applied, leveraging view-averaged gradient magnitude thresholding for point densification, opacity thresholding for pruning, and regular all-points opacity reset. However, we reveal that this strategy is limited in tackling intricate/special image regions (e.g., transparent) as it is unable to identify all the 3D zones that require point densification, and lacking an appropriate mechanism to handle the ill-conditioned points with negative impacts (e.g., occlusion due to false high opacity). To address these limitations, we propose a Localized Point Management (LPM) strategy, capable of identifying those error-contributing zones in the highest demand for both point addition and geometry calibration. Zone identification is achieved by leveraging the underlying multiview geometry constraints, with the guidance of image rendering errors. We apply point densification in the identified zone, whilst resetting the opacity of those points residing in front of these regions so that a new opportunity is created to correct ill-conditioned points. Serving as a versatile plugin, LPM can be seamlessly integrated into existing 3D Gaussian Splatting models. Experimental evaluation across both static 3D and dynamic 4D scenes validate the efficacy of our LPM strategy in boosting a variety of existing 3DGS models both quantitatively and qualitatively. Notably, LPM improves both vanilla 3DGS and SpaceTimeGS to achieve state-of-the-art rendering quality while retaining real-time speeds, outperforming on challenging datasets such as Tanks & Temples and the Neural 3D Video Dataset.
@@ 1 Introduction
Neural rendering has emerged as a generalizable, flexible, and powerful approach for photorealistic novel view synthesis (NVS) of any camera poses [24], underpinning a wide variety of applications in augmented/virtual/mixed reality [9], robotics [39], and generation [26], among more others.
@@
For example, taking a learning-based parametric idea, Neural Radiance Fields (NeRFs) [24] implicitly represent the scene radiance of varying complexity using neural networks (e.g., MLPs), without the tedious requirements of model design handcrafting for accounting the scene variations in geometry, texture, illumination. Despite the high-quality representational performances, they are typically inefficient computationally in view rendering due to heavy ray sampling, thus suffer in scaling to high-resolution content applications and large scale scene modeling [30, 33].
@@
In this context, 3D Gaussian Splatting (3DGS) [13] has come as a more efficient alternative with much faster model optimization and real-time neural rendering. As an explicit representation model, this approach begins by initializing a set of 3D Gaussian points with Structure from Motion (SfM), followed by optimizing the parameters of these points via view reconstruction loss, with the view output produced using a differentiable splatting-based rasterization. As the point initialization is coarse and error-prone, in optimization, a point management mechanism, Adaptive Density Control (ADC), is typically applied for dealing with point distributional issues such as under-population (e.g., no enough points) or over-population (e.g., too many points) in the 3D space. However, we find several limitations with ADC: (1) Thresholding simply the average gradient for deciding the regions for point densification tends to overlook under-optimized points. For example, larger Gaussian points typically exhibit lower average gradients, frequently appearing across various views in screen space. (2) Point sparsity makes it difficult to add sufficient and reliable points to comprehensively cover the scene. (3) Falsely optimized Gaussian points could cause negative effects, e.g., occluding other good points and leading to incorrect depth estimates (see erroneous placements on windows Fig.  1).
@@
To overcome the aforementioned limitations, in this paper we propose a novel Localized Point Management (LPM) approach. Our idea is intuitive – identifying those 3D Gaussian points leading to rendering errors. Thus we start with an image rendering error map of a specific view. To obtain the error contributing 3D points, we leverage the region correspondence between different views via feature mapping, subject to the multiview geometry constraint. For each pair of corresponded regions, we cast the rays through them at their respective camera views in the cone shape, and consider their intersection as the error source zone. Within each such zone, we consider two situations: (1) At presence of points, we further apply point densification at a lower threshold to complement the original counterpart locally; (2) In case no point due to point sparsity, we add new Gaussian points. concurrently, we reset the opacity of points with high estimated opacity and residing in front of these zones, as they could impose high impact on view rendering. This provides an opportunity of correcting those potentially ill-conditioned points whilst tuning those newly added ones in the following optimization. To minimize model expansion, we prune the points by opacity in a density aware manner.
@@
We summarize the contributions as follows: (1) With in-depth analysis, we have discovered that the standard point management mechanism used in 3DGS has several limitations that impede model optimization. (2) We present Localized Point Management (LPM) for these issues by identifying error-contributing 3D zones and implementing appropriate operations for point densification and opacity reset. (3) Extensive experiments validate the benefits of our LPM in improving a diversity of existing 3DGS models in novel view synthesis on both static and dynamic scenes.
@@ 2 Related Work
Neural Scene Representations has always been an important direction in novel view synthesis. Previous methods allocate neural features to structures such as volume [21, 28], texture [32], and point cloud [1]. The pioneering work of NeRF [24] proposes integrating neural networks with 3D volumetric representations to convert a 3D scene into a learnable density field, enabling high-quality novel view synthesis without requiring explicit modeling of the 3D scene and illumination. Later on, numerous works emerge to boost the quality and efficiency of volume rendering, [3, 37, 5] refine the point sampling strategy in ray marching, [4, 35] reparameterize the scene to produce a more compact representation. Additionally, regularization terms [8, 42] can be incorporated to constrain the scene representation, resulting in a closer approximation to real-world geometry. Despite their high-quality representational performance, these methods are typically computationally inefficient for view rendering due to the extensive ray sampling required and the use of Multi-Layer Perceptrons (MLPs) to represent the scene, complicating the computation and optimization of any point within the scene. To address this, several works have proposed novel scene representations aimed at accelerating the rendering process. These representations replace MLPs with sparse voxels [20], hash tables [25], or triplanes [6], significantly enhancing rendering speed. However, real-time rendering remains challenging due to the inherent complexity of the ray marching strategy in volume rendering.
@@
3D Gaussian Splatting represents a recent advancement in novel view synthesis, enabling real-time high-quality rendering. It contributes to splatting-based rasterization by computing pixel colors through depth sorting and $\alpha$-blending of projected 2D Gaussians, thereby avoiding the complex sampling strategies of ray marching and achieving real-time performance. It is precisely due to its real-time high-quality rendering capabilities that 3DGS has been applied to various domains, including autonomous driving, content generation [31], and 4D dynamic scenes [18, 36, 40], among others. Despite these advancements, 3DGS still has some drawbacks, such as the storage of Gaussians and handling multi-resolution, and so on. Several works have enhanced 3DGS by improving Gaussian representation, including techniques such as low-pass filtering [41], multiscale Gaussian representations [38], and interpolating Gaussian attributes from structured grid features [22]. However, these works often overlook the importance of point management, specifically Adaptive Density Control, which is typically applied during optimization to address issues like under-population or over-population in the 3D space. Only a few works have focused on point management. For example, GaussianPro [7] directly tackles densification limitations, bridging gaps from SfM-based initialization. Pixel-GS [43] proposes a gradient scaling strategy to suppress artifacts near the camera. Additionally, [27] introduces an auxiliary per-pixel error function to implicitly supervise point contributions.
@@
Although these methods improve densification, they are still unable to identify all the 3D zones that require point densification and lack a proper mechanism to handle ill-conditioned points with negative impacts. In our paper, we propose a novel approach, Localized Point Management, capable of identifying error-contributing zones with the highest demand for both point addition and geometry calibration.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
Point management Since existing 3DGS variants
start by initializing 3D Gaussian points using Structure from Motion (SfM), the points are often coarse and error-prone. During optimization, a point management mechanism, Adaptive Density Control (ADC), is typically applied to manage point distribution issues.
Specifically, thresholding the average gradient is used to decide on point densification. For each Gaussian point $G_{i}$, 3DGS tracks the magnitude of the positional gradient $\frac{\partial L_{\pi}}{\partial\mu_{i}}$ across all rendered views, which is then averaged to a quantity $T_{i}$. During each training iteration, if the gradient $T_{i}$ surpasses a predefined threshold, it considers this point as inadequately representing the corresponding 3D region. With the scale of the Gaussian as the size measure, a large Gaussian will be split into two, while a small one leads to point cloning.
@@ 3.2 Localized Gaussian Point Management
To address the aforementioned issues, we introduce a novel model agnostic point management approach, Localized Point Management (LPM), which leverages multiview geometry constraints to identify error contributing 3D points, with the guidance of image rendering errors. This approach can be seamlessly integrated with existing 3DGS models without the need for architectural modification.
As illustrated in Figure  2, we begin with an image rendering error map for a specific view. Under the multiview geometry constraint, the corresponding regions in the referred view are matched via feature mapping. For each pair of corresponding regions, we then cast rays through them from their respective camera views in a cone and identify their intersection as the error source zone. Within each zone, we perform localized point manipulation.
@@
Figure 2: Overview of the Localized Point Management (LPM). (a) We start with an image rendering error map against the current view (the ground-truth). Concurrently, matching points are identified between the current view and a refereed view sampled as an adjacent view via off-the-shelf feature mapping. (b)Subsequently, cross-view region mapping is then employed to locate the correspondence region in the refereed view. (c)For each pair of corresponded regions, we cast the rays through them at their respective camera views in the cone shape, and consider their intersection as the error source zone. The final step involves identifying under-optimized or ill-conditioned points within these zones, where under-optimized/empty places are densified, ill-conditioned points are reset.
@@
(i) Cross-view region mapping We select a neighboring view as the referred image. Following LightGlue [19] that predicts a partial assignment between two sets of local features extracted from two view images $A$ and $B$. Each feature consists of sets of 2D features position $\{F_{i}\mid(x_{i},y_{i})\in[0,1]^{2}\}$, normalized by the image size. The images $A$ and $B$ contain $M$ and $N$ local features. LightGlue outputs a set of correspondences =$\mathcal{M}=\{(i,j)\}\subseteq A\times B$. Since the 2D rendering error regions in the current view may not all appear in the referenced image, we select the paired region $(R_{e},R_{e}^{\prime})$ (Figure 2(b)) through the matching points. Additionally, this paired region undergoes multiview adaptive adjustments based on the error map throughout the optimization process.
@@
Recall that in existing 3DGS, points management only relies on the view-averaged gradient magnitude $\tau$ to determine point densification globally. In addition to this, we further perform localized points addition and geometry calibration within the identified error source 3D zone $R_{zone}$. For the point addition, we consider two common situations: (1) In the presence of points, we apply point densification to locally complement the original counterparts. We set a lower threshold to select the points that need densification, aiming to enhance the geometric details. The densification rule is consistent with 3DGS, but it focuses on local 3D zones that need it most. Specifically, for small Gaussians, our strategy involves cloning the Gaussians while maintaining their size and repositioning them along the positional gradient to better capture emerging geometrical features. Conversely, larger Gaussians situated in areas of high variance are split into smaller points to more accurately represent the underlying geometry. (2) In cases of point sparsity, we add new Gaussian points at the center of the 3D zone.
In the context of $\alpha$-blending in 3DGS, if the points at the forefront of the identified 3D zone $R_{zone}$ have the highest opacity, they may occlude valid points, leading to incorrect depth estimation, as shown in Figure 1. To deal with such issues, we treat these points as potentially ill-conditioned. We reset these points to provide an opportunity for correction, further calibrating the geometry.
To minimize model expansion, we adaptively prune points based on their opacity values, starting from low to high. The number of points pruned is determined by the density of points in the zone. This strategic reduction ensures that our point management remains cost efficient and adaptive to the evolving needs of the scene representation.

### v3

# Improving Gaussian Splatting with Localized Points Management

@@ Abstract
Point management is critical for optimizing 3D Gaussian Splatting models, as point initiation (e.g., via structure from motion) is often distributionally inappropriate. Typically, Adaptive Density Control (ADC) algorithm is adopted, leveraging view-averaged gradient magnitude thresholding for point densification, opacity thresholding for pruning, and regular all-points opacity reset. We reveal that this strategy is limited in tackling intricate/special image regions (e.g., transparent) due to inability of identifying all 3D zones requiring point densification, and lacking an appropriate mechanism to handle ill-conditioned points with negative impacts (e.g., occlusion due to false high opacity). To address these limitations, we propose a Localized Point Management (LPM) strategy, capable of identifying those error-contributing zones in greatest need for both point addition and geometry calibration. Zone identification is achieved by leveraging the underlying multiview geometry constraints, subject to image rendering errors. We apply point densification in the identified zones and then reset the opacity of the points in front of these regions, creating a new opportunity to correct poorly conditioned points. Serving as a versatile plugin, LPM can be seamlessly integrated into existing static 3D and dynamic 4D Gaussian Splatting models with minimal additional cost. Experimental evaluations validate the efficacy of our LPM in boosting a variety of existing 3D/4D models both quantitatively and qualitatively. Notably, LPM improves both static 3DGS and dynamic SpaceTimeGS to achieve state-of-the-art rendering quality while retaining real-time speeds, excelling on challenging datasets such as Tanks & Temples and the Neural 3D Video dataset.
@@ 1 Introduction
Neural rendering has emerged as a generalizable, flexible, and powerful approach for photorealistic novel view synthesis (NVS) of any camera poses [23], underpinning a wide variety of applications in augmented/virtual/mixed reality [9], robotics [37], and generation [25], among more others. For example, taking a learning-based parametric idea, Neural Radiance Fields (NeRFs) [23] implicitly represent the scene radiance of any complexity using neural networks (e.g., MLPs), without the tedious requirements of model handcrafting for accounting the scene variations in geometry, texture, illumination. However, their view rendering is inefficient computationally due to heavy ray sampling, thus suffer in scaling to high-resolution content applications and large scale scene modeling [29, 32].
@@
Recently, 3D Gaussian Splatting (3DGS) [14] has come as an alternative with explicit representation, much faster model optimization and real-time neural rendering. The process begins by initializing a set of 3D Gaussian points using Structure from Motion (SfM) [28]. This is followed by optimizing the parameters of these points through view reconstruction loss, resulting in a view output generated with differentiable splatting-based rasterization. However, the point initialization is often distributionally non-optimal, leading to issues such as under-population (e.g., insufficient points) or over-population (e.g., excessive points) in the 3D space. Consequently, a point management mechanism, such as Adaptive Density Control (ADC), is necessary during optimization. However, we identify several limitations with ADC: (i) Thresholding the average gradient to determine regions for point densification often overlooks under-optimized points. For instance, larger Gaussian points typically have lower average gradients and may frequently appear across various views in screen space. (ii) Point sparsity complicates the addition of sufficient and reliable points needed to comprehensively cover the scene. (iii) Mis-optimized Gaussian points can have detrimental effects, such as occluding other valuable points and leading to incorrect depth estimates (see erroneous placements on windows in Fig.  1). While GaussianPro [7] and PixelGS [41] try to solve (1) and (2) through multi-view stereo [2] and additional gradient propagation, respectively, these methods significantly increase the training budget, as shown in the Table  4.
@@
To overcome the aforementioned limitations, in this paper we propose a novel and efficient Localized Point Management (LPM) approach. Our idea is intuitive – identifying those 3D Gaussian points leading to rendering errors. Thus we start with an image rendering error map of a specific view. To obtain the error contributing 3D points, we leverage the region correspondence between different views via feature mapping, subject to the multiview geometry constraint. For each pair of corresponded regions, we cast the rays through them at their respective camera views in the cone shape, and consider their intersection as the error source zone. Within each such zone, we consider two situations: (1) At presence of points, we further apply point densification at a lower threshold to complement the original counterpart locally; (2) In case no point due to point sparsity, we add new Gaussian points. Concurrently, we reset the opacity of points with high opacity estimates that are located in front of these zones, as they can significantly affect view rendering. This provides an opportunity to correct potentially ill-conditioned points while tuning the newly added ones in the subsequent optimization. To minimize model expansion, we prune the points by opacity in a density-aware manner.
@@
We summarize the contributions below: (1) Through in-depth analysis, we have identified several limitations in the standard point management mechanism used in Gaussian Splatting that impede model optimization. (2) We present Localized Point Management (LPM) for these issues by identifying error-contributing 3D zones and implementing appropriate operations for point densification and opacity reset. (3) Extensive experiments validate the benefits of our LPM in improving a diversity of existing 3D and 4D Gaussian Splatting models in novel view synthesis on both static and dynamic scenes.
@@ 2 Related Work
Neural Scene Representations has always been an important direction in novel view synthesis. Previous methods allocate neural features to structures such as volume [21, 27], texture [31], and point cloud [1]. The pioneering work of NeRF [23] proposes integrating neural networks with 3D volumetric representations to convert a 3D scene into a learnable density field, enabling high-quality novel view synthesis without requiring explicit modeling of the 3D scene and illumination. Later on, numerous works emerge to boost the quality and efficiency of volume rendering, [3, 35, 5] refine the point sampling strategy in ray marching, some some advanced works [4, 33] reparameterize the scene to produce a more compact representation. Additionally, regularization terms [8, 39] can be incorporated to constrain the scene representation, resulting in a closer approximation to real-world geometry. Despite their high-quality representational performance, these methods are typically computationally inefficient for view rendering due to the extensive ray sampling required and the use of Multi-Layer Perceptrons (MLPs) to represent the scene, complicating the computation and optimization of any point within the scene. To address this, several works have proposed novel scene representations aimed at accelerating the rendering process. These representations replace MLPs with sparse voxels [20], hash tables [24], or triplanes [6], significantly enhancing rendering speed. However, real-time rendering remains challenging due to the inherent complexity of the ray marching strategy in volume rendering.
@@
Gaussian Splatting represents a recent advancement in novel view synthesis, enabling real-time high-quality rendering. It contributes to splatting-based rasterization by computing pixel colors through depth sorting and $\alpha$-blending of projected 2D Gaussians, thereby avoiding the complex sampling strategies of ray marching and achieving real-time performance. It is precisely due to its real-time high-quality rendering capabilities that 3DGS has been applied to various domains, including autonomous driving, content generation [30], and 4D dynamic scenes [18, 34, 38], among others. Despite these advancements, 3DGS still has some drawbacks, such as the storage of Gaussians and handling multi-resolution, and so on. Several works have enhanced 3DGS by improving Gaussian representation, including techniques such as low-pass filtering [40], multiscale Gaussian representations [36], and interpolating Gaussian attributes from structured grid features [22]. However, these works often overlook the importance of point management, specifically Adaptive Density Control, which is typically applied during optimization to address issues like under-population or over-population in the 3D space. Only a few works have focused on point management. For example, GaussianPro [7] directly tackles densification limitations, bridging gaps from SfM-based initialization. Pixel-GS [41] proposes a gradient scaling strategy to suppress artifacts near the camera. Additionally, [26] introduces an auxiliary per-pixel error function to implicitly supervise point contributions.
@@
Although these methods improve densification, they are still unable to identify all 3D zones that require point densification and lack a proper mechanism to handle ill-conditioned points with negative impacts. Here, we propose a novel approach, Localized Point Management, capable of identifying error-contributing zones with greatest demand for both point addition and geometry calibration.
Figure 2: Overview of our Localized Point Management (LPM). (a) We start with an image rendering error map versus the current view (the ground-truth). Concurrently, matching points are identified between the current view and a referred view sampled as an adjacent view via off-the-shelf feature mapping. (b) Subsequently, cross-view region mapping is then employed to locate the correspondence region in the refereed view. (c) For each pair of corresponded regions, we cast the rays through them at their respective camera views in the cone shape, and consider their intersection as the error source zone. The final step involves identifying under-optimized or ill-conditioned points within these zones, where under-optimized/empty places are densified, and ill-conditioned points are reset.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
Figure 3: Qualitative evaluation of our LPM across diverse static datasets [4, 12]. Our LPM improves 2DGS [13] and 3DGS [14] on these challenging scenarios, e.g. (a) Light artifacts, (b) Completeness in the distance, (c) Depth structure and (d) Mesh details. See red patches for highlighted visual differences.
Point management Since existing 3DGS variants start by initializing 3D Gaussian points using Structure from Motion (SfM), the points are often coarse and non-optimal in space. During optimization, a point management mechanism, Adaptive Density Control (ADC), is typically applied to manage point distribution issues. Specifically, thresholding the average gradient is used to decide on point densification. For each Gaussian point $G_{i}$, 3DGS tracks the magnitude of the positional gradient $\frac{\partial L_{\pi}}{\partial\mu_{i}}$ across all rendered views, which is then averaged to a quantity $T_{i}$. During each training iteration, if the gradient $T_{i}$ surpasses a predefined threshold, it considers this point as inadequately representing the corresponding 3D region. With the scale of the Gaussian as the size measure, a large Gaussian will be split into two, while a small one leads to point cloning.
@@ 3.2 Localized Gaussian Point Management
To address the aforementioned issues, we introduce a novel model agnostic point management approach, Localized Point Management (LPM), which leverages multiview geometry constraints to identify error contributing 3D points, with the guidance of image rendering errors. This approach can be seamlessly integrated with existing 3DGS models without the need for architectural modification. As illustrated in Figure  2, we begin with an image rendering error map for a specific view. Under the multiview geometry constraint, the corresponding regions in the referred view are matched via feature mapping. For each pair of corresponding regions, we then cast rays through them from their respective camera views in a cone and identify their intersection as the error source zone. Within each zone, we perform localized point manipulation.
@@
(i) Cross-view region mapping We select a neighboring view as the referred image. Following LightGlue [19], which predicts a partial assignment between two sets of local features extracted from images $A$ and $B$ from different views. Each feature consists of sets of 2D features position $\{F_{i}\mid(x_{i},y_{i})\in[0,1]^{2}\}$, normalized by the image size. The images $A$ and $B$ contain $M$ and $N$ local features. LightGlue outputs a set of correspondences =$\mathcal{M}=\{(i,j)\}\subseteq A\times B$, where $i$ and $j$ denote the indices of matched points from sets $A$ and $B$, respectively. Since the 2D rendering error regions in the current view may not all appear in the referenced image, we select the paired regions $(R_{e},R_{e}^{\prime})$ (Figure 2(b)) based on the matching points, where $R_{e}$ represents the region in the current view and $R_{e}^{\prime}$ corresponds to the region in the reference view. Additionally, this paired region undergoes multiview adaptive adjustments based on the error map throughout the optimization process.
@@
Method Mip-NeRF 360 Tanks&Temples Deep Blending
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Plenoxels 23.08 0.625 0.463 21.08 0.719 0.379 23.06 0.795 0.510
INGP-Big 25.59 0.699 0.331 21.92 0.745 0.305 24.96 0.817 0.390
Mip-NeRF 360 27.69 0.792 0.237 22.22 0.759 0.257 29.40 0.901 0.245
3DGS 27.21 0.815 0.214 23.14 0.841 0.183 29.41 0.903 0.243
3DGS* 27.47 0.816 0.216 23.67 0.849 0.177 29.55 0.904 0.245
3DGS* + LPM 27.59 0.820 0.216 23.83 0.850 0.181 29.76 0.908 0.241
2DGS* 27.15 0.808 0.246 23.58 0.832 0.185 29.35 0.899 0.262
2DGS* + LPM 27.42 0.817 0.228 23.65 0.848 0.180 29.52 0.903 0.240
MipGS* 27.51 0.817 0.210 23.69 0.852 0.173 29.58 0.910 0.242
MipGS* + LPM 27.70 0.821 0.210 23.82 0.851 0.180 29.61 0.910 0.241
PixelGS* 27.54 0.819 0.203 23.75 0.850 0.175 29.58 0.920 0.220
PixelGS* + LPM 27.80 0.830 0.190 24.02 0.856 0.173 29.65 0.910 0.196
Table 1: Comparison of various methods across different scenes on the Mip-NeRF 360 dataset, Tanks&Temples and Deep Blending. * indicates the retrained model from the official implementation. Bold represents best, underline indicates second best.
@@
Recall that in existing 3DGS, points management only relies on the view-averaged gradient magnitude $\tau$ to determine point densification globally. In addition to this, we further perform localized points addition and geometry calibration within the identified error source 3D zone $R_{zone}$. For the point addition, we consider two common situations: (1) In the presence of points, we apply point densification to locally complement the original counterparts. We set a lower threshold to select the points that need densification, aiming to enhance the geometric details. The densification rule is consistent with 3DGS, but it focuses on local 3D zones that need it most. Specifically, for small Gaussians, our strategy involves cloning the Gaussians while maintaining their size and repositioning them along the positional gradient to better capture emerging geometrical features. Conversely, larger Gaussians situated in areas of high variance are split into smaller points to more accurately represent the underlying geometry. (2) In cases of point sparsity, we add new Gaussian points at the center of the 3D zone. In the context of $\alpha$-blending in 3DGS, if the points at the forefront of the identified 3D zone $R_{zone}$ have the highest opacity, they may occlude valid points, leading to incorrect depth estimation, as shown in Figure 1. To deal with such issues, we treat these points as potentially ill-conditioned. We reset these points to provide an opportunity for correction, further calibrating the geometry. To minimize model expansion, we adaptively prune points based on their opacity values, starting from low to high. The number of points pruned is determined by the density of points in the zone. This strategic reduction ensures that our point management remains cost efficient and adaptive to the evolving needs of the scene representation.